OpenStudy (anonymous):

A model is made of a car. The car is 10 feet long and the model is 7 inches long. What is the ratio of the length of the car to the length of the model? A. 10 : 7 B. 7 : 120 C. 120 : 7 D. 7 : 10 will give medal @phi

OpenStudy (therealmeeeee):

C.120:7

OpenStudy (anonymous):

no im in 11th

OpenStudy (anonymous):

one more question plz

OpenStudy (anonymous):

LHS: small triangle - length = 18 large triangle - length = 27 so ratio = 2:3 (small:big)

OpenStudy (anonymous):

2:3 = 1: 1.5

OpenStudy (anonymous):

so RHS: small triangle - length = 20 large triangle - length = (20 + x) ratio = 2:3 (small:big) so 20 x 1.5 = large triangle length on RHS

OpenStudy (anonymous):

so 20 x 1.5 = (20 + x) 30 = 20 + x ??? = x can u solve from here?

OpenStudy (anonymous):

udue I tried this and I sauck at geometry that's why I came here lol

OpenStudy (anonymous):

nah, it's not geometry, it's logic, so jus learn the rules and when to apply dude and you'll b sweet so 20 x 1.5 = (20 + x) 30 = 20 + x 30-20 = x so x = 10 does this makes sense tho man?

OpenStudy (anonymous):

yes it does thnx a bunch dude

OpenStudy (anonymous):

hall good hey, gud luk man